# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ 5.1 Solution on the Riemann surface
We are now in position to explicitly solve the RHP for the function $W(p)$ (5.12) on the Riemann surface ${\mathcal{R}}$. Define the functions
$w_{1}(p)=\frac{1}{F_{1}(\infty_{1};\tau)}\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{2})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{1})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}e^{-i\delta(p)}\,,$ (5.39)
$w_{2}(p)=\frac{1}{F_{2}(\infty_{2};\tau)}\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{1})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{2})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}e^{-i\delta(p)}\,.$ (5.40)
@@
Note that this determines the zeros $Q_{1},Q_{2}$ of $F_{1}(p;\tau)$ and $F_{2}(p;\tau)$, respectively, although the path of integration from $A$ to $Q_{1}$ or $Q_{2}$ will necessarily involve many traverses of the cycle $\beta$, since $\zeta(x,N)$ is growing in $N$. We can further simplify the expression for $W(p)$ applying the following proposition, which is a Corollary of Lemma 4.1.
@@
$W(p)=\begin{pmatrix}w_{1}(p)\\ w_{2}(p)\end{pmatrix}=\begin{pmatrix}\frac{1}{F_{1}(\infty_{1};\tau)}\frac{\Theta\left(\mathcal{A}(p)+\frac{1+3\tau}{4};\tau\right)\Theta\left(\mathcal{A}(p)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(\mathcal{A}(p)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(p)+\frac{\tau}{2};\tau\right)}e^{-i\delta(p)}\\ \frac{1}{F_{2}(\infty_{2};\tau)}\frac{\Theta\left(\mathcal{A}(p)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(p)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(p)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(p)+\frac{\tau}{2};\tau\right)}e^{-i\delta(p)}\end{pmatrix}\,.$ (5.46)
@@ 5.2 Solution on the Complex plane
First, we explicitly write the norming constant $F_{1}(\infty_{1};\tau),F_{2}(\infty_{2};\tau)$
$F_{1}(\infty):=F_{1}(\infty_{1};\tau)=\frac{\Theta\left(\tau;\tau\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi};\tau\right)}{\Theta\left(\frac{3\tau-1}{4};\tau\right)\Theta\left(\frac{\tau+1}{4};\tau\right)}\,,$ (5.48)
$F_{2}(\infty):=F_{2}(\infty_{2};\tau)=\frac{\Theta\left(0;\tau\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi};\tau\right)}{\Theta\left(\frac{3-\tau}{4};\tau\right)\Theta\left(\frac{1+\tau}{4};\tau\right)}\,,$ (5.49)
then, the entries of the solution (5.47)
$\begin{split}&s_{11}(z;x,t)=\frac{1}{F_{1}(\infty)}\frac{\Theta\left(\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{-i\delta(z)}\\ &s_{12}(z;x,t)=\frac{1}{F_{1}(\infty)}\frac{\Theta\left(-\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{i\delta(z)}\\ &s_{21}(z;x,t)=\frac{1}{F_{2}(\infty)}\frac{\Theta\left(\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{-i\delta(z)}\\ &s_{22}(z;x,t)=\frac{1}{F_{2}(\infty)}\frac{\Theta\left(-\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{i\delta(z)}.\end{split}$ (5.50)
@@ 6 Local parametrices
In the previous section, we developed a solution to RHP 5.1 which serves as an outer approximation to the solution of RHP 4.4 when we are not too close to the four branching points $A,{\overline{A}},-A,-{\overline{A}}$. However, in the vicinity of these points, this approximation is no longer valid. Therefore, we must construct a solution in a small neighborhood around each of these four points to better approximate the solution to RHP 4.4 locally. The global solution to RHP 4.4 will then be given in terms of the error function that describes the difference between the global solution and the approximate solutions, inside and outside the four neighborhoods, which as we will see in this section, exists since it satisfies a small norm RHP.
6.1 Local parametrix near $z={\overline{A}}$
@@
Recall from proposition 4.3 that $\mathcal{U}_{{\overline{A}}}$ is a small neighborhood of $z={\overline{A}}$. Let us denote $E_{{\overline{A}}}$ the solution to the local problem near $z={\overline{A}}$. Below, we outline the process we follow to obtain the solution $E_{{\overline{A}}}$. First, inside $\mathcal{U}_{{\overline{A}}}$, we seek $E_{{\overline{A}}}$ satisfying the Riemann-Hilbert jump relations
@@
$\begin{split}&E_{{\overline{A}}+}(z)=E_{{\overline{A}}-}(z)\,J_{E_{{\overline{A}}}}(z)\\ &J_{E_{{\overline{A}}}}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{4}\\ \\[4.0pt] \mathfrak{N}^{+}(z),\quad&z\in\omega_{1}\cup\omega_{2}\end{cases}\end{split}$ (6.1)
@@
Following [28], we construct $E_{{\overline{A}}}$ in terms of special functions. For this purpose, we define the function $\Psi(\xi)$ as follows
$\Psi(\xi)=\begin{cases}\begin{pmatrix}I_{0}(2\xi^{1/2})&\frac{i}{\pi}K_{0}(2\xi^{1/2})\\ \\[1.0pt] 2\pi i\xi^{1/2}I_{0}^{\prime}(2\xi^{1/2})&-2\xi^{1/2}K_{0}^{\prime}(2\xi^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(-\pi/4,\pi/4)\\ \\[1.0pt] \begin{pmatrix}\frac{1}{2}H_{0}^{(1)}(2(-\xi)^{1/2})&\frac{1}{2}H_{0}^{(2)}(2(-\xi)^{1/2})\\ \\[1.0pt] \pi\xi^{1/2}(H_{0}^{(1)})^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}(H_{0}^{(2)})^{\prime}(2(-\xi)^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(\pi/4,\pi)\\ \\[1.0pt] \begin{pmatrix}\frac{1}{2}H_{0}^{(2)}(2(-\xi)^{1/2})&-\frac{1}{2}H_{0}^{(1)}(2(-\xi)^{1/2})\\ \\[1.0pt] -\pi\xi^{1/2}(H_{0}^{(2)})^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}(H_{0}^{(1)})^{\prime}(2(-\xi)^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(-\pi,-\pi/4)\end{cases}$ (6.2)
@@
where $H_{0}^{(1)}$ and $H_{0}^{(2)}$ are the Hankel functions and $K_{0}$ and $I_{0}$ are the modified Bessel functions, see [12, Chapter 10]. In [28], it is shown that $\Psi$ satisfies the jumps given in Fig. 9. For brevity, we introduce the notation $\mathcal{Y}(z)=2\pi h(z)\rho(z)e^{-2i\theta(z)}$.

### v2

@@ 5.1 Solution on the Riemann surface
Now we can explicitly find $W(p)$ satisfying the conditions (5.12). To this end, we introduce the Abel map related to ${\mathcal{R}}$ as follows
$\mathcal{A}(p)=\int_{A}^{p}\frac{c}{\rho(s)}\,{\rm d}s.$ (5.31)
This is typically multi-valued, and if one wants to create a single-valued function, one may restrict the path of integration to avoid $P_{I}(\gamma)\cup P_{II}(\gamma)\cup P_{I}([{\overline{A}},\infty])\cup P_{II}([{\overline{A}},\infty])$.
Next, we consider the Riemann Theta function [13, 21]
which is an even function of $z$ satisfying the following properties
$\Theta(z+1;\tau)=\Theta(z;\tau)$ (5.33)
$\Theta(z+\tau;\tau)=e^{-\pi i\tau-2\pi iz}\Theta(z;\tau)$ (5.34)
and has a simple zero at $z=\frac{\tau+1}{2}$. We notice that the Riemann Theta function is related to the Jacobi $\theta_{3}$ as $\Theta(z;\tau)=\theta_{3}(\pi z|\tau)$. We use the $\Theta$-function to construct $W(p)$ for $p\in{\mathcal{R}}$, and then through (5.11) we find $\Xi(z)$ solving RHP 5.4. Recall that the function $W(p)$ has simple poles at the points ${\overline{A}},-{\overline{A}}$ and that two entries of $W(p)$ vanish at $\infty_{1}$ and $\infty_{2}$.
Define the functions
$w_{1}(p)=\frac{1}{Z_{1}}\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{2})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{1})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}e^{-i\delta(p)}\,,$ (5.36)
$w_{2}(p)=\frac{1}{Z_{2}}\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{1})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{2})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}e^{-i\delta(p)}\,,$ (5.37)
where
$\begin{split}&Z_{1}=\frac{\Theta\left(\mathcal{A}(\infty_{1})-\mathcal{A}(\infty_{2})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(\infty_{1}))-\mathcal{A}(Q_{1})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(\infty_{1}))-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(\infty_{1}))-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}\\ &Z_{2}=\frac{\Theta\left(\mathcal{A}(\infty_{2})-\mathcal{A}(\infty_{1})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(\infty_{2})-\mathcal{A}(Q_{2})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(\infty_{2})-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(\infty_{2})-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}\,.\end{split}$ (5.38)
@@
Note that this determines the zeros $Q_{1},Q_{2}$ of $w_{1}(p;\tau)$ and $w_{2}(p;\tau)$, respectively, although the path of integration from $A$ to $Q_{1}$ or $Q_{2}$ will necessarily involve many traverses of the cycle $\beta$, since $\zeta(x,N)$ is growing in $N$. We can further simplify the expression for $W(p)$ applying the following proposition, which is a Corollary of (3.22).
@@
$W(p)=\begin{pmatrix}w_{1}(p)\\ w_{2}(p)\end{pmatrix}=\begin{pmatrix}\frac{1}{Z_{1}}\frac{\Theta\left(\mathcal{A}(p)+\frac{1+3\tau}{4};\tau\right)\Theta\left(\mathcal{A}(p)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(\mathcal{A}(p)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(p)+\frac{\tau}{2};\tau\right)}e^{-i\delta(p)}\\ \frac{1}{Z_{2}}\frac{\Theta\left(\mathcal{A}(p)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(p)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(p)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(p)+\frac{\tau}{2};\tau\right)}e^{-i\delta(p)}\end{pmatrix}\,.$ (5.44)
@@ 5.2 Solution on the Complex plane
$\begin{split}&s_{11}(z;x,t)=\frac{1}{Z_{1}}\frac{\Theta\left(\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{-i\delta(z)}\\ &s_{12}(z;x,t)=\frac{1}{Z_{1}}\frac{\Theta\left(-\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{i\delta(z)}\\ &s_{21}(z;x,t)=\frac{1}{Z_{2}}\frac{\Theta\left(\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{-i\delta(z)}\\ &s_{22}(z;x,t)=\frac{1}{Z_{2}}\frac{\Theta\left(-\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}e^{i\delta(z)}.\end{split}$ (5.46)
@@ 6 Local parametrices
In the previous section, we developed a solution to RHP 5.1 which serves as an outer approximation to the solution of RHP 4.4 when we are not too close to the four branching points $A,{\overline{A}},-A,-{\overline{A}}$. However, in the vicinity of these points, this approximation is not expected to be close to the solution of the full Riemann-Hilbert problem (since the jump matrices are not close to $I$ there). Therefore, we must construct a solution in a small neighborhood around each of these four points to better approximate the solution to RHP 4.4 locally. The global solution to RHP 4.4 will then be given in terms of the error function that describes the difference between the global solution and the approximate solutions, inside and outside the four neighborhoods, which as we will see in this section, exists since it satisfies a small norm RHP.
In this section, we construct the local parametrices following [29]. We chose to present the explicit construction of the parametrix near $z=-A$ as it is the most delicate case and requires additional techniques. For the points $z={\overline{A}}$ and $z=A$, we verify that the proposed constructions satisfy the required jump conditions. The local parametrix near $z=-{\overline{A}}$ follows analogously and has therefore been omitted.
6.1 Local parametrix near $z=-A$
@@
Recall from proposition 4.3 that $\mathcal{U}_{-A}$ is a small neighborhood of $z=-A$. Let us denote by $E_{-A}(z)$ a local parametrix near $z=-A$: $E_{-A}(z)$ should satisfy all the Riemann-Hilbert jump relations satisfied by $E(z)$ inside the neighborhood $\mathcal{U}_{-A}$, and it should also behave on the boundary of $\mathcal{U}_{-A}$ so as to asymptotically match the global parametrix.
Inside $\mathcal{U}_{-A}$, we seek $E_{-A}(z)$ satisfying the Riemann-Hilbert jump relations:
@@
$E_{-A+}(z)=E_{-A-}(z)J_{E_{-A}}(z),\quad J_{E_{-A}}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&\quad z\in\eta_{4}\\ \\[2.0pt] e^{\left(-\Delta(x)+\frac{2}{\tau}\ln(N)\right)\sigma_{3}},&\quad z\in\gamma\\ \\[2.0pt] \mathfrak{N}^{+}(z),&\quad z\in\omega_{1}\cup\omega_{2}\end{cases}$ (6.1)
@@
Following [29], we construct $E_{-A}$ in terms of special functions. For this purpose, we define the function $\Psi(\xi)$ as follows:
$\Psi(z)=\begin{cases}\begin{pmatrix}I_{0}(2\xi^{1/2})&\frac{i}{\pi}K_{0}(2\xi^{1/2})\\ \\[1.0pt] 2\pi i\xi^{1/2}I_{0}^{\prime}(2\xi^{1/2})&-2\xi^{1/2}K_{0}^{\prime}(2\xi^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(-2\pi/3,2\pi/3)\\ \\[1.0pt] \begin{pmatrix}\frac{1}{2}H_{0}^{(1)}(2(-\xi)^{1/2})&\frac{1}{2}H_{0}^{(2)}(2(-\xi)^{1/2})\\ \\[1.0pt] \pi\xi^{1/2}(H_{0}^{(1)})^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}(H_{0}^{(2)})^{\prime}(2(-\xi)^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(2\pi/3,\pi)\\ \\[1.0pt] \begin{pmatrix}\frac{1}{2}H_{0}^{(2)}(2(-\xi)^{1/2})&-\frac{1}{2}H_{0}^{(1)}(2(-\xi)^{1/2})\\ \\[1.0pt] -\pi\xi^{1/2}(H_{0}^{(2)})^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}(H_{0}^{(1)})^{\prime}(2(-\xi)^{1/2})\end{pmatrix},\quad&\arg(\xi)\in(-\pi,-2\pi/3)\end{cases}$ (6.2)
@@
where $H_{0}^{(1)}$ and $H_{0}^{(2)}$ are the Hankel functions and $K_{0}$ and $I_{0}$ are the modified Bessel functions, see [13, Chapter 10]. In [29], it is shown that $\Psi$ satisfies the jumps given in Fig. 9. For brevity, we introduce the notation $\mathcal{Y}(z)=2\pi h(z)\rho(z)e^{-2i\theta(z)}$.
@@
We will arrive at precisely this collection of jump relations through explicit transformations, starting with $E_{-A}(z)$.
We introduce the function $m$:
$m(z)=\begin{cases}E_{-A}(z)\,e^{\left(\widetilde{y}(z)-\frac{1}{\tau}\ln(N)\right)\sigma_{3}},\quad&z\in\Omega\\ \\[2.0pt] E_{-A}(z)\,e^{\left(\widetilde{y}(z)+\frac{1}{\tau}\ln(N)\right)\sigma_{3}},\quad&z\in\widetilde{\Omega}.\end{cases}$ (6.3)
$\gamma$$\eta_{4}$$\omega_{2}$$\omega_{1}$$\Omega$$\tilde{\Omega}$
Figure 10: Contour for RHP 6.1
$m_{+}(z)=m_{-}(z)J_{m}(z),$ (6.4)
where the jump matrix $J_{m}(z)$ can be factored as follows
$J_{m}(z)=\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{-\sigma_{3}/2}\left\{\begin{array}[]{lr}\begin{pmatrix}0&i\\ i&0\end{pmatrix},&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{S(s)}}&1\end{pmatrix},&z\in\omega_{1}\cup\omega_{2}\end{array}\right\}\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{\sigma_{3}/2}.$ (6.8)
Note that, in relation (6.8), we have replaced $\int_{-A}^{z}\frac{{\rm d}s}{R(s)}$ with $\int_{-A}^{z}\frac{{\rm d}s}{S(s)}$ since the two agree on the contours $\omega_{1}$ and $\omega_{2}$. It is more convenient to use the function $S(s)$ instead of $R(s)$, because it is branched along $\eta_{3}$ and $\eta_{4}$, which are the relevant contours of non-analyticity in this section.
Next we define the function
$P(z)=m(z)\,\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{-\sigma_{3}/2}\,,$ (6.9)
which satisfies the jump relations
$P_{+}(z)=P_{-}(z)\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{S(s)}}&1\end{pmatrix},\quad&z\in\omega_{1}\cup\omega_{2}.\end{cases}$
Analysis of the local behavior of the function $\frac{4c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{S(s)}$ near $z=-A$ , prompts us to introduce the conformal mapping $\xi=\varphi_{-A}(z)$, with the analytic function $\varphi_{-A}$ defined implicitly via the relation
$\left(\varphi_{-A}(z)\right)^{1/2}\equiv-\ln(N)\frac{c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{S(s)}\,.$ (6.10)
This function maps the point $z=-A$ to $\xi=0$, and the contour $\eta_{4}$ emanating from $-A$ is mapped to $\mathbb{R}_{-}$. We choose the contours $\omega_{1}$ and $\omega_{2}$ (within $\mathcal{U}_{-A}$) so that in the $\xi$ plane, they map to precisely the rays $\mbox{arg}(\xi)=\pm 2\pi/3$ as shown in Figure 9.
Now we consider the jumps in the $\xi$ plane satisfied by $P(\varphi_{-A}^{-1}(\xi))$:
$\left(P(\varphi_{-A}^{-1}(\xi))\right)_{+}=\left(P(\varphi_{-A}^{-1}(\xi))\right)_{-}\begin{cases}\begin{pmatrix}0&-i\\ -i&0\end{pmatrix},\quad&\xi\in(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ ie^{-4\xi^{1/2}}&1\end{pmatrix},\quad&\mbox{Arg}(\xi)=\pm 2\pi/3\end{cases}$ (6.11)
It is then straightforward to verify that the quantity
$P(\varphi_{-A}^{-1}(\xi))e^{2\xi^{1/2}\,\sigma_{3}}e^{\frac{-i\pi\sigma_{3}}{4}}$ (6.12)
satisfies the jump relationships
$\left(P(\varphi_{-A}^{-1}(\xi))e^{2\xi^{1/2}\,\sigma_{3}}e^{\frac{-i\pi\sigma_{3}}{4}}\right)_{+}=\left(P(\varphi_{-A}^{-1}(\xi))e^{2\xi^{1/2}\,\sigma_{3}}e^{\frac{-i\pi\sigma_{3}}{4}}\right)_{-}\begin{cases}\begin{pmatrix}0&1\\ -1&0\end{pmatrix},\quad&\xi\in(-\infty,0)\\ \\[4.0pt] \begin{pmatrix}1&0\\ 1&1\end{pmatrix},\quad&\mbox{Arg}(\xi)=\pm\frac{2\pi}{3},\end{cases}$ (6.13)
which are precisely the jump relationships satisfied by $\Psi$, as shown in Figure 9. We therefore set
$P(\varphi_{-A}^{-1}(\xi))e^{2\xi^{1/2}\,\sigma_{3}}e^{\frac{-i\pi\sigma_{3}}{4}}=\Psi(\xi).$ (6.14)
Then, using the definition (6.2) of $\Psi(\xi)$ and the (conformal and hence invertible) transformation $\varphi_{-A}(z)$, (6.14) defines $P(z)$ back in the $z$ plane. Unraveling the prior transformations from $E_{-A}$ to $m$ and from $m$ to $P$, we have established the following proposition.
Proposition 6.1.
The matrix valued function $E_{-A}$ given by the expression
$E_{-A}(z)=Q_{-A}(z)\,e^{-i\pi\sigma_{3}/4}\,\Psi(\varphi_{-A}(z))\,e^{i\pi\sigma_{3}/4}e^{-2\varphi_{-A}(z)^{1/2}\sigma_{3}}\Bigg{(}\frac{1}{\mathcal{Y}(z)}\Bigg{)}^{\sigma_{3}/2}\begin{cases}e^{-(\widetilde{y}(z)-\frac{1}{\tau}\ln(N))\sigma_{3}},\quad z\in\Omega\\ e^{-(\widetilde{y}(z)+\frac{1}{\tau}\ln(N))\sigma_{3}},\quad z\in\widetilde{\Omega}\end{cases}$ (6.15)
satisfies the jump relation (6.1). Here, $\Psi(\xi)$ is as in (6.2), $Q_{-A}$ is any analytic and invertible matrix-valued function inside $\mathcal{U}_{-A}$ (to be determined later), and the map $\varphi_{-A}$ is defined as
$\varphi_{-A}(z)^{1/2}=-\ln(N)\frac{c}{\tau}\int_{-A}^{z}\frac{{\rm d}s}{S(s)}\,,$ (6.16)
and the domains $\Omega,\,{\widetilde{\Omega}}$ are as in Figure 10.
Remark 6.1.
The factor $e^{-i\pi\sigma_{3}/4}$ in front of the function $\Psi(\varphi_{-A}(z))$ does not arise directly from the construction of $E_{-A}(z)$. However, since the local parametrices in the upper half-plane can be constructed from those in the lower half-plane via a symmetry relation, this factor is included to serve this purpose. Moreover, because it is an analytic function, it does not affect the jump of $E_{-A}(z)$.
6.2 Local parametrix near $z={\overline{A}}$
Recall that $\mathcal{U}_{{\overline{A}}}$ is a small neighborhood of $z={\overline{A}}$. We seek $E_{{\overline{A}}}(z)$, a parametrix which satisfies the same jump relations as $E(z)$ within $\mathcal{U}_{{\overline{A}}}$, and matches the global parametrix on the boundary of $\mathcal{U}_{{\overline{A}}}$. The jump relations for $E$ are:
$E_{+}(z)=E_{-}(z)\,\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{4}\\ \\[4.0pt] \mathfrak{N}^{+}(z),\quad&z\in\omega_{1}\cup\omega_{2}\end{cases}$ (6.17)